# Piecewise M-Stationarity and Related Algorithms for Mathematical Programs with Complementarity Constraints

### 24 Mar 2026 v1

@@ 1.1 First-Order Optimality Condition and Constraint Qualifications
where $\mathcal{T}(\bar{z})$ is the tangent cone of the MPCC at the point $\bar{z}$. If the feasible region is regular at $\bar{z}$ in the sense of Clarke ([25, Definition 6.4][3, Section 1]), this condition is the same as
@@
where $\mathcal{T}(\bar{z})^{*}$ is the dual cone of $\mathcal{T}(\bar{z})$. Without an analytic expression of the tangent cone, verifying these conditions directly is generally nontrivial. In practice, it is desirable to employ linearized cones to reconstruct the first-order condition (4) or (5). Constraint qualifications (CQs) play an important role in this task.
Standard linearization of $\mathcal{T}(\bar{z})$ can be carried out (see [7, Eqs. (10)-(11)]), by replacing the complementarity constraints $0\leq G(z)\perp H(z)\geq 0$ with
@@
and thus justifies using the KKT conditions for RNLP, i.e., the S-stationarity condition, as a necessary first-order condition for the MPCC (see also [8, Theorem 4.1]). We should note that NLP-ACQ cannot hold if $\beta(\bar{z})\neq\emptyset$, because $\mathcal{T}(\bar{z})$ is a nonconvex cone in this case while $\mathcal{T}^{\rm lin}(\bar{z})$ is always a convex polyhedral cone.
@@
then the condition (5), if applicable, can be expressed as $\nabla f(\bar{z})\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$. MPCC-GCQ is implied by MPCC-ACQ, but the converse is in general not true. Their relations are analogous to the relations between NLP-GCQ and NLP-ACQ. Examples showing that NLP-GCQ and MPCC-GCQ have a better chance to be satisfied, even if NLP-ACQ and MPCC-ACQ do not hold, can be found in [28, Example 1.3] and [9, Example 2.1]. Note that despite the fact that a tangent cone is not necessarily equal to the closure of its convex hull, their dual cones are the same. This offers the opportunity for NLP-GCQ and MPCC-GCQ to hold more generally. It has been established that under MPCC-GCQ, M-stationarity is a necessary first-order condition [9, Theorem 3.1]. In addition, Fritz John type M-stationarity has been derived at a local minimizer of an MPCC without requiring a constraint qualification [20, Theorem 3.1].
@@ 1.2 Degeneracy
A difficulty arises in establishing stationarity of a limit point that is degenerate (on the lower level), namely, a sequence $\{z^{k}\}\to\bar{z}$ at which $\beta(\bar{z})\neq\emptyset$. Fukushima and Pang [11] studied the behavior of a sequence $\{z^{k}\}$ that is composed of KKT points of NLPs formulated by smoothing the MPCC with perturbed Fischer-Burmeister functions. The condition of asymptotic weak nondegeneracy was proposed, meaning that for every $i\in\beta(\bar{z})$, $G_{i}(z^{k})$ and $H_{i}(z^{k})$ approach zero at the same order of magnitude. Under this condition and second-order necessary conditions at every $z^{k}$, together with MPCC linear independence constraint qualification (MPCC-LICQ) at $\bar{z}$, it has been proved that $\bar{z}$ is a B-stationary point of the MPCC [11, Theorem 3.1]. However, the condition of asymptotic weak nondegeneracy is hard to enforce in practice. Replacing this condition with upper level strict complementarity (ULSC), namely, $\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}\neq 0$ for all $i\in\beta(\bar{z})$, Scholtes recovered B-stationarity of a limit point of a regularization scheme [27, Corollary 3.4]. Kadrani et al. developed a regularization method whose limit points were shown to be M-stationary under MPCC-LICQ, and S-stationary under additional assumption of asymptotic weak nondegeneracy [18]. The result on M-stationarity was later proved valid under weaker MPCC constant positive linear dependence (MPCC-CPLD) assumption [16]. Results under weaker assumptions also include, for example, that C-stationarity convergence of the method by Steffensen and Ulbrich under MPCC constant rank constraint qualification (MPCC-CRCQ) [29] and under MPCC-CPLD [15], and M-stationarity convergence of the method by Kanzow and Schwartz under MPCC-CPLD [19]. Theoretical and numerical comparison of some of these methods can be found in [16].
@@
Besides diverse methods for reformulating complementarity constraints, many popular algorithmic frameworks in nonlinear programming have been exploited to deal with complementarity as well as the potential degeneracy. The sequential quadratic programming (SQP) methods applied to MPCCs were investigated in [10]. By introducing slack variables into the reformulation of general complementarity constraints, superlinear convergence to a S-stationary point was established under MPCC-LICQ and regularity conditions (Theorems 5.7 and 5.14 therein). An alternative SQP method that retained the superlinear convergence while relaxing some of the assumptions was analyzed in [2], where an adaptive elastic mode was invoked to enforce either feasibility of the QP subproblems or complementarity at the iterates (Theorems 4.5 and 4.6 therein). Interior-penalty methods for MPCCs were studied in [23]; global convergence to a S-stationary point was proved under MPCC-LICQ and a condition on the behavior of the penalty parameters (Theorem 3.4 and Corollary 3.5 therein); superlinear convergence to a S-stationary point was proved under certain regularity conditions (Theorem 4.5 therein); in particular, relations between interior-penalty and interior-relaxation methods were established, which allows to extend some convergence results derived for one approach to the other. Convergence of augmented Lagrangian methods was investigated under MPCC-LICQ [17, Theorem 3.2], where a limit point was proved to be S-stationary in the case of bounded multiplier sequence, and C-stationary in the presence of unbounded multiplier sequences. The results were improved in [1] for a second-order method (Theorem 3.2 therein), where S-stationarity was established under a weaker MPCC-relaxed constant positive linear dependence (MPCC-RCPLD) condition, and convergence in the presence of unbounded multipliers was proved to be M-stationary under MPCC-LICQ. Comparison of some existing augmented Lagrangian methods for MPCCs can be found in [14]. Finally, a detailed survey on these MPCC topics can be found in Kim et al. [21].
@@ 2.1 Equivalence to B-stationarity
On the other hand, since $\bar{z}$ satisfies the first-order necessary condition (8) for optimality, there exists a scalar $\lambda_{0}\geq 0$ and multipliers $\lambda_{I}^{g}\geq 0,\lambda^{h},\lambda_{\alpha}^{G},\lambda_{\gamma}^{H},\zeta$, satisfying the following Fritz John condition ([5, Theorem 6.1.1],[26, Lemma 1 and proof]):
@@
Let $\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z})$ with some $\theta_{i}\in[0,1]$ be an element of the convex hull, then we have (see also [28, Section 2.2])
@@
The MPCC-ACQ assumption in Corollary 2.4 may not be relaxed. To see this, we first note that piecewise M-stationarity of $\bar{z}$ needs the existence of KKT multipliers at $\bar{z}$ for each of the piecewise NLPs. Then, according to [13, Lemma 4.2], we obtain that the condition (12) holds. It follows that $\nabla f(\bar{z})^{T}d\geq 0$ for all $d\in\bigcup_{(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))}\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})=\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})$, whose equivalence to B-stationarity requires the MPCC-ACQ assumption. Furthermore, noting that the sufficient part of this corollary does not rely on MPCC-ACQ, we therefore know that the MPCC-ACQ assumption for the necessary part, i.e., Theorem 2.2, may not be relaxed.
@@ 2.2 Example: Piecewise M-Stationarity Illustration
which relates $\nabla f(\bar{z})$ to the piecewise NLPs. However, we still cannot relate $\nabla f(\bar{z})$ to the MPCC because $\nabla f(\bar{z})\notin\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$, namely,
@@
Comparing (17) and (18), we note that the cones in the latter restrict the choices of $\sigma_{1}$ and $\sigma_{2}$, and therefore prevent (17a) and (17b) holding simultaneously in the case of $\zeta<0$. Choosing $\theta$ according to (16) reveals complementarity of the multipliers $\sigma_{1}$ and $\sigma_{2}$, and therefore piecewise M-stationarity of $\bar{z}$ for the MPCC.
@@ 2.5 Distinction from Extended M-stationarity
  • Since extended M-stationarity is usually difficult to verify in practice [12, Page 924], Definition 4.2 in [12] introduces the concept of strong M-stationarity as a bridge between M-stationarity and S-stationarity. Under an assumption of linearly independent gradients of some MPCC working set of constraints, [12, Theorem 4.3] shows that an extended M-stationary point is strongly M-stationary.
  • A summary of extended M-stationarity and its relation to other properties is given on [12, Page 928] and in [21, Figure 12.2]. Here an extended M-stationary point is also B-stationary, and a B-stationary point that satisfies GGCQ (which we call MPCC-GCQ) is also extended M-stationary.
@@ 3.1 Brief Review of a Bounding Scheme
  • Feasibility: The smoothed NCP function (20) is used to approximate the complementarity constraints in MPCC (1), and the largest difference between them is $\epsilon/2$ [31, Proposition 1.7]. When $\epsilon>0$, every feasible point $z$ of BA($\epsilon$) satisfies
@@
      Suppose that MPCC-LICQ holds at a feasible point of the MPCC. Then in a neighborhood of this point, NLP-LICQ holds at every feasible point of BA($\epsilon$), whenever $\epsilon>0$ is sufficiently small [31, Theorem 3.1].
@@
      Suppose that a sequence of KKT points of problems BA($\epsilon$) tends to a limit point as $\epsilon\to 0$, at which MPCC-LICQ holds. Then the limit point is C-stationary [31, Theorem 3.3].
@@
      In addition, suppose that the reduced Hessian of the Lagrangian at each of the KKT points of problems BA($\epsilon$) is bounded below when $\epsilon>0$ is sufficiently small. Then the limit point is M-stationary [31, Theorem 3.5].
@@ 3.2 Bounding Algorithm
Since $\epsilon/2\leq\Phi_{i}^{\epsilon}(z)\leq 0$ encloses the complementarity condition $0\leq G_{i}(z)\perp H_{i}(z)\geq 0$ [31, Proposition 1.7], we can approximate the complementarity with the equation $\Phi_{i}^{\epsilon}(z)+p_{i}=0$ by letting $p_{i}\in[0,\epsilon/2]$ and $\epsilon\to 0$. For any parameters $p_{i},p_{i}^{\prime}\in[0,\epsilon/2]$ with $\epsilon>0,i=1,\dots,m$, and the corresponding solutions $z(p)$ and $z(p^{\prime})$ to BA($\epsilon$), it is straightforward to show that
@@
This relation explains the estimates of the objective by the Bounding Algorithm. In order to seek a solution to the MPCC, the problem BA($\epsilon^{k}$) is solved to a local solution $z^{k}$, for a sequence of smoothing parameters $\{\epsilon^{k}\}\to 0$. At each solution $z^{k}=z(p^{k})$, we take advantage of the sensitivities $u^{\Phi,k}=u^{\Phi}(p^{k})$ to adjust the parameters $p^{k}$, with the aim of improving the objective in the subsequent solution. As shown in Step 3 of the algorithm, in the case of $u_{i}^{\Phi,k}>0$ (or $u_{i}^{\Phi,k}<0$), increasing (or decreasing) $p_{i}^{k}$ may lead to decrease in the objective function, and the corresponding indices make up a set $P_{0}$ (or $P_{\epsilon}$). Step 4 then adjusts the parameters for the next solution; $p_{i}^{k+1}$ takes the extreme values for $i\in P_{0}\cup P_{\epsilon}$; otherwise, $p_{i}^{k+1}$ is set to $\kappa p_{i}^{k}$ for $i\notin P_{0}\cup P_{\epsilon}$, where the constant $\kappa\in(0,1)$. When $\epsilon^{k}>0$ is sufficiently small, $z(p^{k})$ is an $\epsilon^{k}$-approximate solution to the MPCC, which includes an $O((\epsilon^{k})^{2})$ correction arising from the parameters adjustment [31, Proposition 2.1].

### 14 Apr 2026 v2

@@ 1.1 First-Order Optimality Condition and Constraint Qualifications
where $\mathcal{T}(\bar{z})$ is the tangent cone of the MPCC at the point $\bar{z}$. This condition is the same as
@@
where $\mathcal{T}(\bar{z})^{*}$ is the dual cone of $\mathcal{T}(\bar{z})$ (see also [24, Theorem 6.12]). Without an analytic expression of the tangent cone, verifying these conditions directly is generally nontrivial. In practice, it is desirable to employ linearized cones to reconstruct the first-order condition (4) or (5). Constraint qualifications (CQs) play an important role in this task.
Standard linearization of $\mathcal{T}(\bar{z})$ can be carried out (see [6, Eqs. (10)-(11)]), by replacing the complementarity constraints $0\leq G(z)\perp H(z)\geq 0$ with
@@
and thus justifies using the KKT conditions for RNLP, i.e., the S-stationarity condition, as a necessary first-order condition for the MPCC (see also [7, Theorem 4.1]). We should note that NLP-ACQ cannot hold if $\beta(\bar{z})\neq\emptyset$, because $\mathcal{T}(\bar{z})$ is a nonconvex cone in this case while $\mathcal{T}^{\rm lin}(\bar{z})$ is always a convex polyhedral cone.
@@
then the condition (5) can be expressed as $\nabla f(\bar{z})\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$. MPCC-GCQ is implied by MPCC-ACQ, but the converse is in general not true. Their relations are analogous to the relations between NLP-GCQ and NLP-ACQ. Examples showing that NLP-GCQ and MPCC-GCQ have a better chance to be satisfied, even if NLP-ACQ and MPCC-ACQ do not hold, can be found in [27, Example 1.3] and [8, Example 2.1]. Note that despite the fact that a tangent cone is not necessarily equal to the closure of its convex hull, their dual cones are the same. This offers the opportunity for NLP-GCQ and MPCC-GCQ to hold more generally. It has been established that under MPCC-GCQ, M-stationarity is a necessary first-order condition [8, Theorem 3.1]. In addition, Fritz John type M-stationarity has been derived at a local minimizer of an MPCC without requiring a constraint qualification [19, Theorem 3.1].
@@ 1.2 Degeneracy
A difficulty arises in establishing stationarity of a limit point that is degenerate (on the lower level), namely, a sequence $\{z^{k}\}\to\bar{z}$ at which $\beta(\bar{z})\neq\emptyset$. Fukushima and Pang [10] studied the behavior of a sequence $\{z^{k}\}$ that is composed of KKT points of NLPs formulated by smoothing the MPCC with perturbed Fischer-Burmeister functions. The condition of asymptotic weak nondegeneracy was proposed, meaning that for every $i\in\beta(\bar{z})$, $G_{i}(z^{k})$ and $H_{i}(z^{k})$ approach zero at the same order of magnitude. Under this condition and second-order necessary conditions at every $z^{k}$, together with MPCC linear independence constraint qualification (MPCC-LICQ) at $\bar{z}$, it has been proved that $\bar{z}$ is a B-stationary point of the MPCC [10, Theorem 3.1]. However, the condition of asymptotic weak nondegeneracy is hard to enforce in practice. Replacing this condition with upper level strict complementarity (ULSC), namely, $\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}\neq 0$ for all $i\in\beta(\bar{z})$, Scholtes recovered B-stationarity of a limit point of a regularization scheme [26, Corollary 3.4]. Kadrani et al. developed a regularization method whose limit points were shown to be M-stationary under MPCC-LICQ, and S-stationary under additional assumption of asymptotic weak nondegeneracy [17]. The result on M-stationarity was later proved valid under weaker MPCC constant positive linear dependence (MPCC-CPLD) assumption [15]. Results under weaker assumptions also include, for example, that C-stationarity convergence of the method by Steffensen and Ulbrich under MPCC constant rank constraint qualification (MPCC-CRCQ) [28] and under MPCC-CPLD [14], and M-stationarity convergence of the method by Kanzow and Schwartz under MPCC-CPLD [18]. Theoretical and numerical comparison of some of these methods can be found in [15].
@@
Besides diverse methods for reformulating complementarity constraints, many popular algorithmic frameworks in nonlinear programming have been exploited to deal with complementarity as well as the potential degeneracy. The sequential quadratic programming (SQP) methods applied to MPCCs were investigated in [9]. By introducing slack variables into the reformulation of general complementarity constraints, superlinear convergence to a S-stationary point was established under MPCC-LICQ and regularity conditions (Theorems 5.7 and 5.14 therein). An alternative SQP method that retained the superlinear convergence while relaxing some of the assumptions was analyzed in [2], where an adaptive elastic mode was invoked to enforce either feasibility of the QP subproblems or complementarity at the iterates (Theorems 4.5 and 4.6 therein). Interior-penalty methods for MPCCs were studied in [22]; global convergence to a S-stationary point was proved under MPCC-LICQ and a condition on the behavior of the penalty parameters (Theorem 3.4 and Corollary 3.5 therein); superlinear convergence to a S-stationary point was proved under certain regularity conditions (Theorem 4.5 therein); in particular, relations between interior-penalty and interior-relaxation methods were established, which allows to extend some convergence results derived for one approach to the other. Convergence of augmented Lagrangian methods was investigated under MPCC-LICQ [16, Theorem 3.2], where a limit point was proved to be S-stationary in the case of bounded multiplier sequence, and C-stationary in the presence of unbounded multiplier sequences. The results were improved in [1] for a second-order method (Theorem 3.2 therein), where S-stationarity was established under a weaker MPCC-relaxed constant positive linear dependence (MPCC-RCPLD) condition, and convergence in the presence of unbounded multipliers was proved to be M-stationary under MPCC-LICQ. Comparison of some existing augmented Lagrangian methods for MPCCs can be found in [13]. Finally, a detailed survey on these MPCC topics can be found in Kim et al. [20].
@@ 2.1 Equivalence to B-stationarity
On the other hand, since $\bar{z}$ satisfies the first-order necessary condition (8) for optimality, there exists a scalar $\lambda_{0}\geq 0$ and multipliers $\lambda_{I}^{g}\geq 0,\lambda^{h},\lambda_{\alpha}^{G},\lambda_{\gamma}^{H},\zeta$, satisfying the following Fritz John condition ([4, Theorem 6.1.1],[25, Lemma 1 and proof]):
@@
Let $\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z})$ with some $\theta_{i}\in[0,1]$ be an element of the convex hull, then we have (see also [27, Section 2.2])
@@
The MPCC-ACQ assumption in Corollary 2.4 may not be relaxed. To see this, we first note that piecewise M-stationarity of $\bar{z}$ needs the existence of KKT multipliers at $\bar{z}$ for each of the piecewise NLPs. Then, according to [12, Lemma 4.2], we obtain that the condition (12) holds. It follows that $\nabla f(\bar{z})^{T}d\geq 0$ for all $d\in\bigcup_{(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))}\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})=\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})$, whose equivalence to B-stationarity requires the MPCC-ACQ assumption. Furthermore, noting that the sufficient part of this corollary does not rely on MPCC-ACQ, we therefore know that the MPCC-ACQ assumption for the necessary part, i.e., Theorem 2.2, may not be relaxed.
@@ 2.2 Example: Piecewise M-Stationarity Illustration
which relates $\nabla f(\bar{z})$ to the piecewise NLPs. However, we still cannot relate $\nabla f(\bar{z})$ to the MPCC because
@@
Comparing (17) and (18), we note that the cones in the latter restrict the choices of $\sigma_{1}$ and $\sigma_{2}$, and therefore prevent (17a) and (17b) holding simultaneously in the case of $\zeta<0$. Choosing $\theta$ according to (16) reveals complementarity of the multipliers $\sigma_{1}$ and $\sigma_{2}$ (i.e., $\sigma_{1}\sigma_{2}=0$), and therefore piecewise M-stationarity of $\bar{z}$ for the MPCC.
@@ 2.5 Distinction from Extended M-stationarity
  • Since extended M-stationarity is usually difficult to verify in practice [11, Page 924], Definition 4.2 in [11] introduces the concept of strong M-stationarity as a bridge between M-stationarity and S-stationarity. Under an assumption of linearly independent gradients of some MPCC working set of constraints, [11, Theorem 4.3] shows that an extended M-stationary point is strongly M-stationary.
  • A summary of extended M-stationarity and its relation to other properties is given on [11, Page 928] and in [20, Figure 12.2]. Here an extended M-stationary point is also B-stationary, and a B-stationary point that satisfies GGCQ (which we call MPCC-GCQ) is also extended M-stationary.
@@ 3.1 Brief Review of a Bounding Scheme
  • Feasibility: The smoothed NCP function (20) is used to approximate the complementarity constraints in MPCC (1), and the largest difference between them is $\epsilon/2$ [30, Proposition 1.7]. When $\epsilon>0$, every feasible point $z$ of BA($\epsilon$) satisfies
@@
      Suppose that MPCC-LICQ holds at a feasible point of the MPCC. Then in a neighborhood of this point, NLP-LICQ holds at every feasible point of BA($\epsilon$), whenever $\epsilon>0$ is sufficiently small [30, Theorem 3.1].
@@
      Suppose that a sequence of KKT points of problems BA($\epsilon$) tends to a limit point as $\epsilon\to 0$, at which MPCC-LICQ holds. Then the limit point is C-stationary [30, Theorem 3.3].
@@
      In addition, suppose that the reduced Hessian of the Lagrangian at each of the KKT points of problems BA($\epsilon$) is bounded below when $\epsilon>0$ is sufficiently small. Then the limit point is M-stationary [30, Theorem 3.5].
@@ 3.2 Bounding Algorithm
Since $\epsilon/2\leq\Phi_{i}^{\epsilon}(z)\leq 0$ encloses the complementarity condition $0\leq G_{i}(z)\perp H_{i}(z)\geq 0$ [30, Proposition 1.7], we can approximate the complementarity with the equation $\Phi_{i}^{\epsilon}(z)+p_{i}=0$ by letting $p_{i}\in[0,\epsilon/2]$ and $\epsilon\to 0$. For any parameters $p_{i},p_{i}^{\prime}\in[0,\epsilon/2]$ with $\epsilon>0,i=1,\dots,m$, and the corresponding solutions $z(p)$ and $z(p^{\prime})$ to BA($\epsilon$), it is straightforward to show that
@@
This relation explains the estimates of the objective by the Bounding Algorithm. In order to seek a solution to the MPCC, the problem BA($\epsilon^{k}$) is solved to a local solution $z^{k}$, for a sequence of smoothing parameters $\{\epsilon^{k}\}\to 0$. At each solution $z^{k}=z(p^{k})$, we take advantage of the sensitivities $u^{\Phi,k}=u^{\Phi}(p^{k})$ to adjust the parameters $p^{k}$, with the aim of improving the objective in the subsequent solution. As shown in Step 3 of the algorithm, in the case of $u_{i}^{\Phi,k}>0$ (or $u_{i}^{\Phi,k}<0$), increasing (or decreasing) $p_{i}^{k}$ may lead to decrease in the objective function, and the corresponding indices make up a set $P_{0}$ (or $P_{\epsilon}$). Step 4 then adjusts the parameters for the next solution; $p_{i}^{k+1}$ takes the extreme values for $i\in P_{0}\cup P_{\epsilon}$; otherwise, $p_{i}^{k+1}$ is set to $\kappa p_{i}^{k}$ for $i\notin P_{0}\cup P_{\epsilon}$, where the constant $\kappa\in(0,1)$. When $\epsilon^{k}>0$ is sufficiently small, $z(p^{k})$ is an $\epsilon^{k}$-approximate solution to the MPCC, which includes an $O((\epsilon^{k})^{2})$ correction arising from the parameters adjustment [30, Proposition 2.1].